A biologist has found the average weight of 12 randomly selected mud turtles to be 8.7 lb with standard deviation 3.6 lb. Find a 90% confidence interval for the population mean weight of all such turtles.
Answer by stanbon(48569)

A biologist has found the average weight of 12 randomly selected mud turtles to be 8.7 lb with standard deviation 3.6 lb. Find a 90% confidence interval for the population mean weight of all such turtles.
--------------
E = 1.645*3.6/sqrt(12) = 1.5829....
C.I. = (8.7-1.5829, 8.7+1.5829)
====================
Cheers,
Stan H.